OpenStudy (anonymous):

Square root of 4 million?

OpenStudy (anonymous):

2000

OpenStudy (anonymous):

How do you find it though? Tried and didnt work

OpenStudy (anonymous):

idk

OpenStudy (amistre64):

sqrt(4) and half the zeros

OpenStudy (anonymous):

i knew it

OpenStudy (anonymous):

= (4 x 1 000 000)^(1/2) = (4)^(1/2) x (10^6)^(1/2) = 2 x 10^3 = 2000